# Adversarial Diffusion for Robust Reinforcement Learning

### v1

@@ 3.1 Markov Decision Processes and Reinforcement Learning
Consider a Markov Decision Process (MDP) $M=\langle S,A,P,r,\gamma,\rho\rangle$, where $S$ and $A$ are the state and action spaces, respectively, $P(\cdot|s,a)$ is the transition probability function, $r(\cdot|s,a)$ is the reward function, $\gamma$ is the discount factor and $\rho$ is the initial state distribution. By interacting with the MDP, a Reinforcement Learning agent is able to collect sequences of states, actions and rewards, forming trajectories $\boldsymbol{\tau}=(s_{0},a_{0},r_{0},\dots,s_{H},a_{H},r_{H})$. The objective of the Reinforcement Learning agent is to learn an optimal policy $\pi^{\star}$ maximizing the policy value $v_{\pi}(s)=\mathbb{E}_{\pi}[\sum_{i=0}^{\infty}\gamma^{i}r_{t+i+1}|s_{t}=s]$. In this paper, we consider a model-based RL setting, where we use a diffusion model to approximate the distribution of trajectories under a given policy. Specifically, if $p^{\pi}$ denotes the true distribution of the trajectories $\boldsymbol{\tau}$ under policy $\pi$, the diffusion model samples trajectories with distribution $p_{\boldsymbol{\theta}}$ close to $p^{\pi}$. We adopt a Dyna-style approach sutton1991dyna , where the diffusion model and the policy are iteratively updated: the policy is improved using data collected from the model, while the model is improved using samples gathered from the target environment using the learned policy.
@@ 5 Algorithms
We now introduce Adversarial Diffusion for Robust Reinforcement Learning (AD-RRL), which alternates between model improvement and policy improvement steps (see Algorithm 1). AD-RRL leverages the adversarial conditional sampling discussed in the previous section to sample trajectories in the worst $\alpha$-percentile in terms of return. Our approach is primarily inspired by PolyGRAD rigter2023world and Diffuser janner2022planning .
@@ 6 Experiments
In Figure˜1, we plot the return under the different algorithms and for selected environments and varying parameters. Additional plots are provided in Section˜F.2. In most environments, AD-RRL consistently outperforms both robust and non-robust baselines. PPO and TRPO appear surprisingly stable, which is likely a consequence of the well-tuned Stable-Baselines3 implementations—but are still matched or surpassed by AD-RRL. At the same time, AD-RRL consistently outperforms both DR-PPO and M2TD3, demonstrating greater stability and achieving higher cumulative rewards across all environments.
Furthermore, a direct comparison with PolyGRAD (the foundation of our algorithm) highlights that our modifications significantly improve performance under diverse test-time conditions, enhancing robustness to large parameter shifts and model misspecifications. This can be clearly seen in Figure˜1(d) or Figure˜1(g). In the Reacher environment (Figure˜1(i)) the difference in performance is less evident, but our model still performs consistently better or on par with the baselines. It is also clear from Figures˜1(a) and 1(b) that while PolyGRAD achieves slightly better performance on the nominal environment (as observed in Table˜1), it sacrifices robustness under perturbed conditions.
For some environments—see for example Figure˜1(a) and Figure 3(b) (presented in Appendix F), AD-RRL performance degrades for extreme changes in the modified parameter (but it remains better than other algorithms). We hypothesize that this is because our model is generating challenging trajectories which are nonetheless plausible under the agent policy and environment dynamics. Extreme changes in the environment physics do not reflect these constraints, and relevant trajectories might not be generated often.
@@
Sample efficiency. The learning curves in Section˜F.1 (see Appendix F) show that AD-RRL reaches higher or matching final performance with the same number of samples as the baselines, and in several cases converges faster. Hence, our adversarially guided diffusion not only preserves (or improves) performance on the training environment (with nominal physics parameters), but also matches the sample efficiency of state-of-the-art model-based and model-free alternatives.
@@ 8 Acknowledgments
When comparing the numerators of both terms in Equation˜22, since $\log\eta_{i}(\alpha,N)=\frac{1}{N}\log\left(\frac{1}{\alpha}\right)$, for $N$ large enough, $\sqrt{2\log\eta_{i}(\alpha,N)}<R-||\mu_{\boldsymbol{\theta}}(\boldsymbol{\tau}_{i},i)||_{\infty}$.
@@ Appendix E Implementation details
When computing $c_{i}$ according to (18), we chose $R=3\sigma_{i}$, where $\sigma_{i}$ is the standard deviation of the diffusion process [^2] In Denoising Diffusion Probabilistic Models, the standard deviation is fixed at every step $i$ according to a known scheduling rate [15]. at step $i$. In our implementation we choose $\eta_{i}(\alpha,N)=\left(\frac{1}{\alpha}\right)^{\frac{1}{N}}$.
@@ F.1 Training curves
Figure˜2 shows the learning curves for AD-RRL and all the baselines for the considered MuJoCo tasks. Across seeds, AD-RRL reaches its final performance at least as quickly as the other methods. AD-RRL also achieves a final score matching or surpassing that of the baselines. This is particularly clear for the Cheetah and Walker environments, presented in Figures˜2(d) and 2(e). These results confirm that our method is more robust to modeling errors but does not sacrifice optimality in the training environment or learning speed.
@@
Figure 2: Training-return curves for five MuJoCo tasks. Shaded areas represent one standard error.
@@ F.2 Varying parameters
In Figure˜3 we present additional parameters variations for the InvertedPendulum and Cartpole environment. The pattern is consistent with the plots presented in Figure˜1: AD-RRL achieves on par or higher returns than both robust and non-robust baselines as the dynamics deviate from nominal values. The only exception appears to be for higher variations of the cart mass, in the Inverted Pendulum environment (Figure˜3(b)), where the additional inertia pushes most methods toward failure and AD-RRL similarly shows a performance decline.
@@
Figure 3: Average Return for varying physical parameters. Shaded areas represent one standard error.

### v2

@@ 1 Introduction
Figure 1: A high-level overview of AD-RRL. Following a Dyna-like structure sutton1991dyna , the algorithm iteratively: samples trajectories from the real environment with a given policy $\pi$, improves the Diffusion models on the collected data, uses an Adversarially Guided Diffusion model (Section 4) to generate challenging synthetic trajectories which in turn are used to improve $\pi$. The loop is repeated until convergence to the optimal policy $\pi^{\star}$.
@@ 3.1 Markov Decision Processes and Reinforcement Learning
Consider a Markov Decision Process (MDP) $M=\langle S,A,P,r,\gamma\rangle$, where $S$ and $A$ are the state and action spaces, respectively, $P(\cdot|s,a)$ is the transition probability function, $r(\cdot|s,a)$ is the reward function and $\gamma$ is the discount factor. By interacting with the MDP, a Reinforcement Learning agent is able to collect sequences of states, actions and rewards, forming trajectories $\boldsymbol{\tau}=(s_{0},a_{0},r_{0},\dots,s_{H},a_{H},r_{H})$. The objective of the Reinforcement Learning agent is to learn an optimal policy $\pi^{\star}$ maximizing the policy value $v_{\pi}(s)=\mathbb{E}_{\pi}[\sum_{i=0}^{\infty}\gamma^{i}r_{t+i+1}|s_{t}=s]$. In this paper, we consider a model-based RL setting, where we use a diffusion model to approximate the distribution of trajectories under a given policy. Specifically, if $p^{\pi}$ denotes the true distribution of the trajectories $\boldsymbol{\tau}$ under policy $\pi$, the diffusion model samples trajectories with distribution $p_{\boldsymbol{\theta}}$ close to $p^{\pi}$. We adopt a Dyna-style approach sutton1991dyna , where the diffusion model and the policy are iteratively updated: the policy is improved using data collected from the model, while the model is improved using samples gathered from the target environment using the learned policy.
@@ 5 Algorithms
We now introduce Adversarial Diffusion for Robust Reinforcement Learning (AD-RRL) [^2] The official implementation for AD-RRL can be found on GitHub: https://github.com/danielefoffano/AD-RRL, which alternates between model improvement and policy improvement steps (see Algorithm 1). AD-RRL leverages the adversarial conditional sampling discussed in the previous section to sample trajectories in the worst $\alpha$-percentile in terms of return. Our approach is primarily inspired by PolyGRAD rigter2023world and Diffuser janner2022planning .
@@ 6 Experiments
In Figure˜2, we plot the return under the different algorithms and for selected environments and varying parameters. Additional plots are provided in Section˜F.2. In most environments, AD-RRL consistently outperforms both robust and non-robust baselines. PPO and TRPO appear surprisingly stable, which is likely a consequence of the well-tuned Stable-Baselines3 implementations—but are still matched or surpassed by AD-RRL. At the same time, AD-RRL consistently outperforms both DR-PPO and M2TD3, demonstrating greater stability and achieving higher cumulative rewards across all environments.
Furthermore, a direct comparison with PolyGRAD (the foundation of our algorithm) highlights that our modifications significantly improve performance under diverse test-time conditions, enhancing robustness to large parameter shifts and model misspecifications. This can be clearly seen in Figure˜2(d) or Figure˜2(g). In the Reacher environment (Figure˜2(i)) the difference in performance is less evident, but our model still performs consistently better or on par with the baselines. It is also clear from Figures˜2(a) and 2(b) that while PolyGRAD achieves slightly better performance on the nominal environment (as observed in Table˜1), it sacrifices robustness under perturbed conditions.
For some environments—see for example Figure˜2(a) and Figure 4(b) (presented in Appendix F), AD-RRL performance degrades for extreme changes in the modified parameter (but it remains better than other algorithms). We hypothesize that this is because our model is generating challenging trajectories which are nonetheless plausible under the agent policy and environment dynamics. Extreme changes in the environment physics do not reflect these constraints, and relevant trajectories might not be generated often.
@@
Sample efficiency. The learning curves in Figure˜3 (see Appendix F.1) show that AD-RRL reaches higher or matching final performance with the same number of samples as the baselines, and in several cases converges faster. Hence, our adversarially guided diffusion not only preserves (or improves) performance on the training environment (with nominal physics parameters), but also matches the sample efficiency of state-of-the-art model-based and model-free alternatives. In Section˜F.3, we perform an ablation study comparing AD-RRL trained on 1.51.5M samples against the Model-Free Baselines trained on 33M samples. This way we allow the model-free methods to compensate for their lower sample efficiency by leveraging their computational efficiency. Overall, Figures˜5 and 6 show that even when model-free baselines are given twice as many samples, AD-RRL remains on average better or on par across the proposed tasks.
@@ 8 Acknowledgments
When comparing the numerators of both terms in Equation˜22, since $\log\eta_{i}(\alpha,N)=\frac{1}{N}\log\left(\frac{1}{\alpha}\right)$, for $N$ large enough, $\sqrt{2\log\eta_{i}(\alpha,N)}<\rho-||\mu_{\boldsymbol{\theta}}(\boldsymbol{\tau}_{i},i)||_{\infty}$.
@@ Appendix E Implementation details
When computing $c_{i}$ according to (18), we chose $\rho=3\sigma_{i}$, where $\sigma_{i}$ is the standard deviation of the diffusion process [^3] In Denoising Diffusion Probabilistic Models, the standard deviation is fixed at every step $i$ according to a known scheduling rate [15]. at step $i$. In our implementation we choose $\eta_{i}(\alpha,N)=\left(\frac{1}{\alpha}\right)^{\frac{1}{N}}$.
@@ F.1 Training curves
Figure˜3 shows the learning curves for AD-RRL and all the baselines for the considered MuJoCo tasks. Across seeds, AD-RRL reaches its final performance at least as quickly as the other methods. AD-RRL also achieves a final score matching or surpassing that of the baselines. This is particularly clear for the Cheetah and Walker environments, presented in Figures˜3(d) and 3(e). These results confirm that our method is more robust to modeling errors but does not sacrifice optimality in the training environment or learning speed.
@@
Figure 3: Training-return curves on the nominal environment for five MuJoCo tasks. Shaded areas represent one standard error over 5 runs.
@@ F.2 Varying parameters
In Figure˜4 we present additional parameters variations for the InvertedPendulum and Cartpole environment. The pattern is consistent with the plots presented in Figure˜2: AD-RRL achieves on par or higher returns than both robust and non-robust baselines as the dynamics deviate from nominal values. The only exception appears to be for higher variations of the cart mass, in the Inverted Pendulum environment (Figure˜4(b)), where the additional inertia pushes most methods toward failure and AD-RRL similarly shows a performance decline.
@@
Figure 4: Average Return for varying physical parameters. Shaded areas represent one standard error over 5 runs.
F.3 Ablation study
In this section, we compare AD-RRL trained on 1.51.5M environment samples with model-free baselines trained on 33M samples. This setup allows model-free methods to compensate for their lower sample efficiency by leveraging their greater computational efficiency. In Figure˜5 we present the learning curves on the nominal environment for the Model-Free algorithms. We can observe that AD-RRL still outperforms or (in the worst case) matches the performance of the Model-Free baselines in all of the environments. This is also reflected when testing the robustness of the algorithms, as shown in Figure˜6. AD-RRL still achieves on par or higher returns than robust and non-robust Model-Free baselines. When comparing with Figure˜2, we can see that some baselines perform worse than their 1.51.5M samples versions (e.g., DR-PPO on the Walker and Hopper environments; TRPO on Hopper). This is expected, since an increased amount of samples might improve the performance on the nominal training environment, but will also lead to overfitting, making the agents more vulnerable to parameter changes in the test environment. The only exception appears to be M2TD3, which is able to outperform AD-RRL on the Walker environment at test time, when varying the physics parameters, despite achieving the same performance in the nominal environment. However, M2TD3 seems to be more prone to overfitting in easier environments, as seen in Figure˜5(a), where the baseline cannot reach the optimal solution but instead suffers a substantial performance decline in the latest training stages. Finally, M2TD3 remains stuck to a suboptimal policy in the Hopper environment, similarly to what we observed when training it on 1.51.5M samples. The doubled amount of samples does not seem to benefit the learning process of this baseline. Overall, even when model-free baselines are given twice as many samples, AD-RRL remains on average better or on par across the proposed tasks.
(a) InvertedPendulum
(b) Reacher
(c) Hopper
(d) Cheetah
(e) Walker
Figure 5: Training-return curves on the nominal environment for the Model-Free baslines trained on 33M samples. Shaded areas represent one standard error over 5 runs. The dashed line represents the reference final cumulative reward achieved by AD-RRL trained on 1.51.5M samples.
(a) Hopper – Mass
(b) Hopper – Gravity
(c) InvertedPendulum - Mass Pole
(d) InvertedPendulum – Mass Cart
(e) Hopper – Friction
(f) Cheetah - Gravity
(g) Cheetah – Mass
(h) Cheetah – Friction
(i) Reacher – Actuator Gear
(j) Walker – Friction
(k) Walker – Mass
(l) InvertedPendulum – Gravity
Figure 6: Average return across variations in selected physics parameters. Shaded regions indicate $\pm$ one standard error over 5 runs. AD-RRL and PolyGRAD are trained on 1.51.5M samples, while the remaining Model-Free Baselines are trained on 33M samples.